# Autoregressive Processes on Stiefel and Grassmann Manifolds

### v1

@@ Abstract
keywords:
Statistics on manifolds , differential geometry , Homogeneous spaces
2020 MSC:
62R30 , 53Z50
journal: JOURNALTOFIND
[…]
[upps]organization=Department of Mathematics, Uppsala University,addressline=Ångströmslaboratoriet, Lägerhyddsvägen 1, city=Uppsala, postcode=PO Box 480, 751 06, country=Sweden
@@ 1 Introduction
Now consider a sequence of measurements over time, a time-series that takes values on a Grassmann or Stiefel manifold. If one now wants to make a prediction of future states or analyse the dynamics of such a time series, we have to make a suitable model for the time series. In $\mathbb{R}$-valued statistics, the most basic model used in time series analysis are the autoregressive processes. This then begs the question: What is a natural definition of an autoregressive process on Stiefel and Grassmann manifolds? Moreover, how can one estimate the parameters of the model in a reasonable manner?
@@
In [7], Xavier and Manton introduced some AR-processes on Riemannian manifolds, a treatise we draw inspiration from and aim to specialise to the class of homogeneous spaces. The goal is to enable us to estimate the system parameters on Grassmann and Stiefel manifolds. Since, by definition, a homogeneous space has a transitive Lie-group action, this naturally guides us to consider system parameters as elements in the Lie group. In Proposition 3.1, we show that one may model an AR($N$)-processes on Stiefel and Grassmann manifolds by considering the system parameter as an element in the orthogonal group, see Equation (2). Moreover, by Theorem 3.6 we show that the system parameter can be estimated, and numerically computed, by averaging out over the observations. Using the definition for the expectation on manifolds, we propose a conjugate gradient descent method in Algorithm 1, which utilise an approximation of the distance function for the Grassmann and Stiefel manifolds.
@@
This paper is relevant to anyone doing directional statistics, the most common form is that of directional statistics on the spheres $\mathbb{S}^{n}$. The spheres are special cases of the Stiefel manifolds, more precisely $\mathbb{S}^{n}\cong\operatorname{St}_{n+1,1}$. Ever since Fischer in [8] popularized the idea of modelling data on spheres in order to ascertain geological information on rocks, see also [9], there has been an abundance of applications of directional statistics. Directional statistics has found its way into meteorology [10, 11, 12, 13], into medicine [14, 15] and into machine learning [16], to name a few.
@@
We start, in Section 2, by giving an overview on what has been done previously in the area of autoregressive processes on Riemannian manifolds, and general statistics on manifolds. Then, in Section 3, we specialise to homogeneous spaces and we give some new results. Thereafter, in Section 4, we give some algorithms that estimate and approximate the system parameters on the Stiefel and Grassmann manifold using the conjugate gradient method. A fundamental introduction to Riemannian geometry is given in Appendix A that introduces concepts needed in order to understand the reasoning in this paper. In the second part of the appendix, Appendix B, one can find some figures showing simulations for a select few Stiefel and Grassmann manifolds illustrating the algorithms given in Section 4.
@@ 2 Statistics on Riemannian Manifolds
Throughout, let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space. Let $M$ be a Riemannian manifold. We shall recall a slightly modified, but still equivalent, version of the ”wrapped process” definition of an AR-process found in [7]. Denote $\exp_{p}$ and $\log_{p}$ to be the Riemannian exponential map centred at $p$ and its inverse, respectively, see Definition A.2. Moreover, the parallel transport from $p\in M$ to $q\in M$ along a geodesic from $p$ to $q$ is written as $P_{p,q}:T_{p}M\rightarrow T_{q}M$.
@@
where $\varepsilon_{\ell}:\Omega\rightarrow T_{Z_{\ell-1}}M$ is a random variable with $\mathbb{E}[\varepsilon_{\ell}]=0$, $\operatorname{Var}(\varepsilon_{\ell})=\operatorname{id}_{T_{Z_{\ell-1}}M}$.
For our purposes we need the notion of expected value of a Riemann manifold-valued random variable. This is precisely the well known least square miminiser on $\mathbb{R}^{n}$ adapted and generalized to Riemannian manifolds. Suppose $\operatorname{dist}$ is the Riemannian distance function on $M$ and $\operatorname{Vol}$ is the Riemannian volume measure on $M$. Following [17, 18], the expectation of an absolutely continuous random variable $X:(\Omega,\mathcal{F},\mathbb{P})\rightarrow M$ is then any point $y\in M$ satisfying
$\mathbb{E}[X]=\operatorname*{arginf}_{y\in M}\int_{M}\operatorname{dist}^{2}(y,x)p_{X}(x)\operatorname{dVol}(x).$
@@ 3 AR-Processes on Homogeneous Spaces by Lie Group Action
$Z_{\ell}=\exp_{Z_{\ell-1}}\left(\varepsilon_{\ell}+\sum_{j=1}^{N}P_{Z_{\ell-j},Z_{\ell-1}}\log_{Z_{\ell-j}}(\Phi_{j}Z_{\ell-j})\right).$
@@
be an absolutely continuous random variable. Suppose $m=\mathbb{E}[X]$ exists and is unique. Assume $p_{X}(y)$ is reflection symmetric around $m$. Let $f:M\rightarrow M$ be an injective, nonexhaustive and reflection symmetric function and assume its inverse $f^{-1}:M\rightarrow M$ is reflection symmetric as well. Then, the probability density function of $f(X)$ is reflection symmetric around $f(m)$.
@@
In particular the (Matrix) Lie-group action is reflection symmetric as the following lemma illustrates.
@@
Let $G$ be a matrix Lie-group with bi-invariant Riemannian metric, and let $M$ be a $G$-homogeneous space with induced Riemannian metric from $G$. Then, for any $p,q\in M$, and any $\Phi\in G$, the following holds:
  1. 1.
    $\operatorname{dist}(\Phi p,\Phi q)\leq\operatorname{dist}(p,q);$
  2. 2.
    Let $X=\log_{p}(q)$, and $\hat{q}=\exp(-X)p$, then
@@
$i)$: Let $X\in T_{e}G$ be horizontal to $M$ such that $q=\exp_{\mathrm{M}}(X)p$ then $\operatorname{dist}(p,q)=\norm{X}$. Now it holds that
$\Phi q$ $=\Phi\exp_{\mathrm{M}}(X)p$
$=\Phi\exp_{\mathrm{M}}(X)\Phi^{-1}\Phi p$
$=\exp_{\mathrm{M}}(\Phi X\Phi^{-1})\Phi p.$
But $\Phi X\Phi^{-1}$ is not necessarily horizontal and thus
$\operatorname{dist}(\Phi p,\Phi q)=\norm{\pi_{\text{Horizontal}}\left(\Phi X\Phi^{-1}\right)}_{G}\leq\norm{\Phi X\Phi^{-1}}_{G}=\norm{X}_{G}=\operatorname{dist}(p,q),$
where $\pi_{\text{Horizontal}}$ denotes the projection into the horizontal direction for $M$.
ii)ii): Suppose $X\in T_{e}G$ is horizontal such that $q=\exp_{\mathrm{M}}(X)p$ and set $\hat{q}=\exp_{\mathrm{M}}(-X)p$. Then, it holds similarly to $i)$ that
@@
In other words Lemma 3.4 shows that the Lie group action is Lipschitz with Lipschitz constant $1$, i.e. non-expansive, and it shows that it is reflection symmetric around all points. Note also that when $M$ is a $G$-symmetric space we have equality in $i)$. Before we go to the main theorem we shall need to check that the space of reflection symmetric variables on Euclidean-space indeed is closed under addition.
@@
Let
@@
be an AR(1)-process. Suppose that $\varepsilon_{\ell}$, $Z_{\ell-1}$ are absolutely continuous w.r.t. $\mathbb{P}$ and $\mathbb{E}[\varepsilon_{\ell}]=0$. Moreover, assume that $\varepsilon_{\ell}$ and $Z_{\ell-1}$ are reflection symmetric around $0\in T_{Z_{\ell}}M$ and $\mathbb{E}[Z_{\ell-1}]$, respectively, and that they are all mutually independent for all $\ell$. Then, it holds that
@@
Set $x=\mathbb{E}[Z_{\ell-1}]$, then we need to show that the function $f:M\rightarrow[0,\infty)$ defined by
@@
is minimal precisely when $y=\Phi\mathbb{E}[Z_{\ell-1}]$. To see this, consider the gradient of $f$ w.r.t. $y$. By definition of the distance function, it holds that $\nabla_{y}\operatorname{dist}(y,z)^{2}=2\log_{y}$. Therefore,
$\nabla f(y)=\int_{M}p_{Z_{\ell}}(z)2\log_{y}(z)\operatorname{dVol}_{M}(z)$
which will be zero at $y=\Phi x$ if $p_{Z_{\ell}}$ is rotationally symmetric around $\Phi x$ since $\log_{\Phi x}(z)$ is anti-symmetric around $\Phi x$. Moreover,
$\nabla^{2}f(y)\mid_{y=\Phi x}=2I_{T_{y}M},$
which is positive definite.
Left to show is that $p_{Z_{\ell}}(z)$ indeed is reflection symmetric around $\Phi x$ from that $\varepsilon_{\ell}$ and $Z_{\ell-1}$ are reflection symmetric. By Lemma 3.4 the Lie group action satisfies the assumptions required for $f$ in Lemma 3.3 and therefore $p_{\Phi Z_{\ell-1}}$ is reflection symmetric with mean $\Phi x$. Moreover, by construction, the Riemannian logarithm is a local isometry, and thus we may once again apply Lemma 3.3 to $\log_{Z_{\ell-1}}(\Phi Z_{\ell-1})$. Next we have sum of two independent refection symmetric random variables is reflection symmetric by Lemma 3.5 and hence
$\varepsilon_{\ell}+\log_{Z_{\ell-1}}(\Phi Z_{\ell-1})$
is reflection symmetric around $\log_{x}(\Phi x)$. Lastly $\exp_{x}$ is a local isometry again and by Lemma 3.3 $Z_{\ell-1}$ is reflection symmetric with expectation
$\exp_{x}(\log_{x}(\Phi x))=\Phi x=\Phi\mathbb{E}[Z_{\ell-1}]$
This follows from that the Riemannian exponential and logarithm are by construction local isometries, Lemma 3.4 ii)ii) and a similar symmetry argument, this time on $Z_{\ell-1}$, the probability density function of $Z_{\ell}$ is necessarily centred at $\Phi x$. ∎
@@ 4.1 Orthogonal/Unitary groups
for every $\ell$. Then the average $\hat{\Phi}$ may be computed using a barycentre algorithm on all the $\hat{\Phi}_{\ell}$ as in [20]. The numerics turn out to work as expected as seen in Figure 1.
@@ 4.2 Stiefel Manifolds
Consider $M=\operatorname{St}_{n,k}$, the real compact Stiefel manifold with the natural left $O(n)$-action. We want to miminise the equation $f:O(n)\rightarrow\mathbb{R}_{\geq 0}$ defined by
@@ 4.3 Grassmann Manifolds
Now since $X^{T}\Delta=A$ is antisymmetric we have the following Grassmann metric at the base point $X\in\operatorname{St}_{n,k}$
@@
note the difference of a factor of $\frac{1}{2}$ compared to the Stiefel case. Moreover, note that this metric is degenerate on the Stiefel manifold but non-degenerate precisely on the Grassmann manifold. Therefore the map we want to miminise can be approximated as follows:
@@
A Basic Riemannian Geometry

### v2

@@ Abstract
Keywords Statistics on manifolds; differential geometry; Homogeneous spaces; directional statistics
2020 Mathematics subject classification 62R30; 53Z50
@@ 1 Introduction
Now consider a sequence of measurements over time, a time-series that takes values on a Grassmann or Stiefel manifold. If one now wants to make a prediction of future states or analyze the dynamics of such a time series, we have to make a suitable model for the time series. In $\mathbb{R}$-valued statistics, the most basic model used in time series analysis are the autoregressive processes. This then begs the question: What is a natural definition of an autoregressive process on Stiefel and Grassmann manifolds? Moreover, how can one estimate the parameters of the model in a reasonable manner?
@@
In [23], Xavier and Manton introduced some AR-processes on Riemannian manifolds, a treatise we draw inspiration from and aim to specialize to the class of homogeneous spaces. The goal is to enable us to estimate the system parameters on Grassmann and Stiefel manifolds. Since, by definition, a homogeneous space has a transitive Lie-group action, this naturally guides us to consider system parameters as elements in the Lie group. In Proposition 3.1, we show that one may model an AR($N$)-processes on Stiefel and Grassmann manifolds by considering the system parameter as an element in the orthogonal group, see Equation (4). Moreover, by Theorem 3.6 we show that the system parameter can be estimated, and numerically computed, by averaging out over the observations. Using the definition for the expectation on manifolds, we propose a conjugate gradient descent method in Algorithm 1, which utilize an approximation of the distance function for the Grassmann and Stiefel manifolds.
@@
This paper is relevant to anyone doing directional statistics, the most common form is that of directional statistics on the spheres $\mathbb{S}^{n}$. The spheres are special cases of the Stiefel manifolds, more precisely $\mathbb{S}^{n}\cong\operatorname{St}_{n+1,1}$. Ever since Fischer in [6] popularized the idea of modeling data on spheres in order to ascertain geological information on rocks, see also [16], there has been an abundance of applications of directional statistics. Directional statistics has found its way into meteorology [1, 14, 17, 18], into medicine [2, 9] and into machine learning [8], to name a few.
@@
We start, in Section 2, by giving an overview on what has been done previously in the area of autoregressive processes on Riemannian manifolds, and general statistics on manifolds. Then, in Section 3, we specialize to homogeneous spaces and we give some new results. Thereafter, in Section 4, we give some algorithms that estimate and approximate the system parameters on the Stiefel and Grassmann manifold using the conjugate gradient method. A fundamental introduction to Riemannian geometry is given in Appendix A that introduces concepts needed in order to understand the reasoning in this paper. In the second part of the appendix, Appendix B, one can find some figures showing simulations for a select few Stiefel and Grassmann manifolds illustrating the algorithms given in Section 4.
@@ 2 Statistics on Riemannian Manifolds
Throughout, let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space. Let $M$ be a Riemannian manifold. We shall recall a slightly modified, but still equivalent, version of the ”wrapped process” definition of an AR-process found in [23]. Denote $\exp_{p}$ and $\log_{p}$ to be the Riemannian exponential map centered at $p$ and its inverse, respectively, see Definition A.2. Moreover, the parallel transport from $p\in M$ to $q\in M$ along a geodesic from $p$ to $q$ is written as $P_{p,q}:T_{p}M\rightarrow T_{q}M$.
@@
where $\varepsilon_{\ell}:\Omega\rightarrow T_{Z_{\ell-1}}M$ is a random variable with $\mathbb{E}[\varepsilon_{\ell}]=0$, $\operatorname{Cov}(\varepsilon_{\ell})=\operatorname{id}_{T_{Z_{\ell-1}}M}$.
For our purposes we need the notion of expected value of a Riemann manifold-valued random variable. This is precisely the well known least square minimizer on $\mathbb{R}^{n}$ adapted and generalized to Riemannian manifolds. Suppose $\operatorname{dist}$ is the Riemannian distance function on $M$ and $\operatorname{Vol}$ is the Riemannian volume measure on $M$. Following [3, 19], the expectation of an absolutely continuous random variable $X:(\Omega,\mathcal{F},\mathbb{P})\rightarrow M$ is then any point $y\in M$ satisfying
$\mathbb{E}[X]=\operatorname*{arginf}_{y\in M}\int_{M}\operatorname{dist}^{2}(y,x)p_{X}(x)\operatorname{dVol}(x).$ (1)
@@ 3 AR-Processes on Homogeneous Spaces by Lie Group Action
$Z_{\ell}=\exp_{Z_{\ell-1}}\left(\varepsilon_{\ell}+\sum_{j=1}^{N}P_{Z_{\ell-j},Z_{\ell-1}}\log_{Z_{\ell-j}}(\Phi_{j}Z_{\ell-j})\right)$ (3)
and with $\operatorname{Cov}(\varepsilon_{\ell})=\sigma^{2}\operatorname{id}_{T_{Z_{\ell-1}}M}$.
@@
be an absolutely continuous random variable. Suppose $m=\mathbb{E}[X]$ exists and is unique. Assume $p_{X}(y)$ is reflection symmetric around $m$. Let $f:M\rightarrow M$ be an injective, non-exhaustive and reflection symmetric function and assume its inverse $f^{-1}:M\rightarrow M$ is reflection symmetric as well. Then, the probability density function of $f(X)$ is reflection symmetric around $f(m)$.
@@
It is well known that for any if $M$ is a Riemannian $G$-homogeneous space and if then, action from $G$ acts as isometries, that is it holds that
$\operatorname{dist}(\Phi p,\Phi q)=\operatorname{dist}(p,q)$ (5)
for all $p,q\in M$ and all $\Phi\in G$. In other words (5) shows that the Lie group action is Lipschitz with Lipschitz constant $1$, i.e. non-expansive, and we shall see that (Matrix) Lie group action is reflection symmetric as well.
@@
Let $G$ be a matrix Lie-group with bi-invariant Riemannian metric, and let $M$ be a $G$-homogeneous space with induced Riemannian metric from $G$. Then, for any $p,q\in M$, and any $\Phi\in G$, it holds that if $X=\log_{p}(q)$, and $\hat{q}=\exp(-X)p$, then
@@
Suppose $X\in T_{e}G$ is horizontal such that $q=\exp_{\mathrm{M}}(X)p$ and set $\hat{q}=\exp_{\mathrm{M}}(-X)p$. Then, it holds that
@@
Before we go to the main theorem we shall need to check that the space of reflection symmetric variables on Euclidean-space indeed is closed under addition.
@@
Let $M$ be a Riemannian $G$-homogeneous space with $G$ a compact matrix Lie-group and let
@@
be an AR(1)-process on $M$ as in (3). Suppose that $\varepsilon_{\ell}$ and $Z_{\ell-1}$ are absolutely continuous w.r.t. $\mathbb{P}$ with respective unique expected values $\mathbb{E}[\varepsilon_{\ell}]=0$ and $\mathbb{E}[Z_{\ell-1}]$. Moreover, assume that $\varepsilon_{\ell}$ and $Z_{\ell-1}$ are reflection symmetric around $0\in T_{Z_{\ell}}M$ and $\mathbb{E}[Z_{\ell-1}]$, respectively, and that they are all mutually independent for all $\ell$. Furthermore, suppose $\rho$ is the injectivity radius of $M$ and assume that $\varepsilon_{\ell}$ is such that $\mathbb{P}(\norm{\varepsilon_{\ell}}\leq\rho/2)=1$. Then, it holds that $Z_{\ell}$ is reflection symmetric around $\Phi\mathbb{E}[Z_{\ell-1}]$ and
@@
Set $x=\mathbb{E}[Z_{\ell-1}]$. First we show that $p_{Z_{\ell}}(z)$ is reflection symmetric around $\Phi x$ from that $\varepsilon_{\ell}$ and $Z_{\ell-1}$ are reflection symmetric. By Lemma 3.4 the Lie group action satisfies the assumptions required for $f$ in Lemma 3.3 and therefore $p_{\Phi Z_{\ell-1}}$ is reflection symmetric around $\Phi x$. Moreover, the Riemannian logarithm map is a local isometry, and thus we may once again apply Lemma 3.3 to $\log_{Z_{\ell-1}}(\Phi Z_{\ell-1})$. Next we have sum of two independent refection symmetric random variables is reflection symmetric by Lemma 3.5 and hence
$\varepsilon_{\ell}+\log_{Z_{\ell-1}}(\Phi Z_{\ell-1})$
is reflection symmetric around $\log_{x}(\Phi x)$. Lastly $\exp_{x}$ is a local isometry again and by Lemma 3.3 $Z_{\ell-1}$ is reflection symmetric around
$\exp_{x}(\log_{x}(\Phi x))=\Phi x=\Phi\mathbb{E}[Z_{\ell-1}]$
This follows from that the Riemannian exponential and logarithm are by construction local isometries, Lemma 3.4 ii)ii) and a similar symmetry argument, this time on $Z_{\ell-1}$, the probability density function of $Z_{\ell}$ is necessarily centered at $\Phi x$.
Next we need to show that the function $f:M\rightarrow[0,\infty)$ defined by
@@
is minimal precisely when $y=\Phi x$. To see this, consider the gradient of $f$ w.r.t. $y$. By definition of the distance function, it holds that $\nabla_{y}\operatorname{dist}(y,z)^{2}=2\log_{y}$. Therefore,
$\nabla f(y)=-\int_{M}p_{Z_{\ell}}(z)2\log_{y}(z)\operatorname{dVol}_{M}(z)$
which will be zero at $y=\Phi x$ since $p_{Z_{\ell}}$ is rotationally symmetric around $\Phi x$ and since $\log_{\Phi x}(z)$ is anti-symmetric around $\Phi x$. Note that
$\mathbb{E}[Z_{\ell}\mid Z_{\ell-1}=u]=\operatorname*{arginf}_{m\in M}\int_{M}\operatorname{dist}(m,z)^{2}p_{Z_{\ell}\mid Z_{\ell-1}=u}(z)\operatorname{dVol}(z)$
and by definition it holds that
$p_{Z_{\ell}\mid Z_{\ell-1}=u}(z)=p_{\exp_{u}(\varepsilon_{\ell}+\log_{u}(\Phi u))}(z).$
Since $\varepsilon_{\ell}$ is reflection symmetric around $0\in T_{u}M$ and since $\varepsilon_{\ell}$ is independent of the event $Z_{\ell-1}=u$, it holds that $p_{Z_{\ell}\mid Z_{\ell-1}}$ is reflection symmetric around $\Phi u$. This follows from observing that Lemma 3.5 applies to $\varepsilon_{\ell}+\log_{u}(\Phi u)$ and that Proposition 3.3 applies to $\exp_{u}(\varepsilon_{\ell}+\log_{u}(\Phi u))$ since $\exp_{u}$ is a local isometry. Now since $M$ is a Riemannian homogeneous space of a compact Lie group, $M$ is compact and it has bounded sectional curvature, therefore [10, Theorem 1.2] applies using that $\mathbb{P}(\norm{\varepsilon_{\ell}}\leq\rho/2)=1$ and we can conclude that
$\mathbb{E}[Z_{\ell}\mid Z_{\ell-1}=u]=\operatorname*{arginf}_{m\in M}\int_{M}\operatorname{dist}(\Phi u,z)^{2}p_{Z_{\ell}\mid Z_{\ell-1}=u}(z)\operatorname{dVol}(z)=\Phi u.$
Hence, it holds that
$\mathbb{E}[Z_{\ell}]=\mathbb{E}[\mathbb{E}[Z_{\ell}\mid Z_{\ell-1}=u]]=\mathbb{E}[\Phi Z_{\ell-1}].$
Now by a change of variables by the map $h:M\rightarrow M$, defined by $h(u)=\Phi u$, with inverse $h^{-1}(u)=\Phi^{-1}u$, it holds that
[…]
$\mathbb{E}[Z_{\ell}]=\Phi\mathbb{E}[Z_{\ell-1}].$
@@ 4.1 Orthogonal/Unitary groups
for every $\ell$. Then the average $\hat{\Phi}$ may be computed using a barycenter algorithm on all the $\hat{\Phi}_{\ell}$ as in [20]. The numerics turn out to work as expected as seen in Figure 1.
@@ 4.2 Stiefel Manifolds
Consider $M=\operatorname{St}_{n,k}$, the real compact Stiefel manifold with the natural left $O(n)$-action. We want to minimize the equation $f:O(n)\rightarrow\mathbb{R}_{\geq 0}$ defined by
@@ 4.3 Grassmann Manifolds
Now since $X^{T}\Delta=A$ is anti-symmetric we have the following Grassmann metric at the base point $X\in\operatorname{St}_{n,k}$
@@
note the difference of a factor of $\frac{1}{2}$ compared to the Stiefel case. Moreover, note that this metric is degenerate on the Stiefel manifold but non-degenerate precisely on the Grassmann manifold. Therefore the map we want to minimize can be approximated as follows:
@@
Appendix A Basic Riemannian Geometry